12. The demand for April, May and June are 130, 128 and 137 respectively. Compute a weighted three-month moving average where the weights are highest for the latest months and descend in the order of 3. 2 and 1.
1
Expert's answer
2019-10-21T12:37:38-0400

Weighted three-month average as per weights


"WMA={\sum Weight \times Data\ Value \over \sum Weight}"

"WMA={ 3 \times 130+2\times 128+1\times 137 \over 1+2+3}=130.5\ units"